A boat, which moves at 20 miles per hour in water without a current, goes 198 miles upstream and 198 miles back again in 20 hours. Find the speed of the current to the nearest tenth.

Answer by mananth(16946)

Boat speed speed = 20 mph
current speed = x mph
0.60
against current = 20 - x
with current = 20 + x

Total distance = 198 miles Total Time= 20.00 hours
Time upstream + time downstream =

198 /( 20 - x )+ 198 /( 20 + x )= 20
198 ( 20 + x )+ 198 ( 20 - x )= 20.00 ( 400 - 1 X^2 )
3960 + 198 x + 3960 -198 x = 8000 - 20 X^2
7920 - -8000 = -20 X^2
-80 = -20 X^2
-20 X^2= -80
X^2 = 4
X = +/- 2
it can't be negative